OpenStudy (anonymous):

Find the points on the lemniscate where the tangent is horizontal. 2(x^2 + y^2)^2 = 25(x^2 − y^2)

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

Familiar with implicit differentiation?

OpenStudy (anonymous):

I know I have to use it I just struggle with it.

OpenStudy (anonymous):

ok, well the plan is to use implicit differentiation to solve for dy/dx, then solve for (x,y) : dy/dx = 0.

OpenStudy (anonymous):

Remember that implicit differentiation means using the chain rule. What is being implied is that y =f(x), so any derivative of y must include y'

OpenStudy (anonymous):

i need 4 points....

zepdrix (zepdrix):

Did you find your derivative yet toots? :o

OpenStudy (anonymous):

yes sir!

OpenStudy (anonymous):

\[4(x^2+y^2)(2x+2y*y')=25(x^2-y^2)(2x-2y*y')\]

zepdrix (zepdrix):

The initial expression, was the right side supposed to be squared as well?

OpenStudy (anonymous):

no lol

zepdrix (zepdrix):

Ok then we wouldn't have this,\[4(x^2+y^2)(2x+2y*y')=25\cancel{(x^2-y^2)}(2x-2y*y')\]

zepdrix (zepdrix):

You took the derivative of that, that's what the other brackets are.

OpenStudy (anonymous):

\[25(2x-2y*y')\]

zepdrix (zepdrix):

Hmmm I'm trying to decide whether we should solve for y' or just plug in y'=0 from here.

OpenStudy (anonymous):

you are the boss! :)

zepdrix (zepdrix):

Ya let's just plug in y'=0 from here. I'm thinking it will work out nicely :o

zepdrix (zepdrix):

y' represents the slope of the lines tangent to the curve. When y'=0 the slope of the tangent is zero (horizontal).

zepdrix (zepdrix):

So we need to find the locations at which that is happening.

OpenStudy (anonymous):

\[4(x^2+y^2)(2x)=25(2x)\]

OpenStudy (anonymous):

i did not even see you put that equation up lol

zepdrix (zepdrix):

XD my bad, stealing all the fun. I'll erase it hehe

OpenStudy (anonymous):

you didn't have to lol i was happy i got the same thing

OpenStudy (anonymous):

now do i solve for y?

OpenStudy (usukidoll):